According to the expectations hypothesis, What do you expect the interest rate on a two-year bond to be three years from now? The current interest rates for different maturities are as follows:

Maturity Rate
1-year 3.00
2-year 3.50
3-year 3.75
4-year 4.00
5-year 4.00

Enter your answer as a percent without the % sign. Round your final answer to two decimals.
Answer #1

(1+F2)^2 = [1+i5]^5 /[1+i3]^3

(1+F2)^2 = [1+.04]^5 /[1+ .0375]^3

(1+F2)^2 = [1.04]^5/[1.0375]^3

   (1+F2)^2 = 1.21665/ 1.11677

(1+F2)^2 = 1.0894

   Multiplying both sides by ^1/2

(1+F2)^1/2*2 = (1.0894)^1/2

(1+F2) = 1.0438

F2 = 1.0438 -1

              = .0438 OR 4.38%